# Toward Robust Imperceptible Perturbation against Unauthorized Text-to-image Diffusion-based Synthesis

### v2

@@ 7 Conlusion
This paper proposes MetaCloak, the first work that protects user images from unauthorized subject-driven text-to-image synthesis under data transformation defense. MetaCloak resolves the limitations of existing works in sub-optimal optimization and fragility to data transformations with a novel meta-learning framework and transformation-robust perturbation crafting process. Extensive experiments demonstrate that MetaCloak can effectively degrade the diffusion-based personalized generation under various training settings with and without data transformations. Notably, MetaCloak is practical and can be applied to protect images in a black-box manner against online training-as-service platforms like Replicate. An important future direction is to establish a theoretical foundation for the effectiveness and robustness of MetaCloak. Another interesting direction is to design a more efficient and robust perturbations training approach.
@@ A.1 Hardware and DreamBooth Training Details
All the experiments are conducted on an Ubuntu 20.04.6 LTS (focal) environment with 503GB RAM, 10 GPUs (NVIDIA® RTX® A5000 24GB) and 32 CPU cores (Intel® Xeon® Silver 4314 CPU @ 2.40GHz). Python 3.10.12 and Pytorch 1.13 are used for all the implementations. For the DreamBooth full trianing mode, we use the 8-bit Adam optimizer [Kingma and Ba, 2017] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ under bfloat16-mixed precision and enable the xformers for memory-efficient training. For calculating prior loss, we use 200 images generated from Stable Diffusion v2-1-base with the class prompt “a photo of a person”. The weight for prior loss is set to 1. For the instance prompt, we use “a photo of sks person”. During the meta-learning process, we regularly delete the temporary models and store the surrogates back to the CPU to save GPU memory. It takes about 3 GPU hours to craft perturbations for an instance under this strategy.
@@ B.1 Evaluation Metrics
In this section, we describe the evaluation metrics used in our experiments in more detail. For our proposed CLIP-IQAC, we calculate the CLIP score difference between “a good photo of [class]” and “a bad photo of [class]”. For calculating SDS and IMS-VGGNet, we leverage the APIs for face recognition and face embedding extraction in the deep face library Serengil and Ozpinar [2021]. In terms of graphical quality, we found that the commonly used metric, BRISQUE Mittal et al. [2012] is not a faithful metric when we conduct additional data transformations like Gaussian Filtering. We thus omit this score when presenting results in the main text. For instance, the BRISQUE score of a fully poisoned Dreambooth is better than the clean one, as shown in Tab. 7. Among all the considered metrics, we found that SDS and IMS-VGG are more aligned with our perception of evaluating the personalized generation performance of Dreambooth. The SDS score indicates whether a subject is presented in the generated image, while the IMS-VGGNet score measures the similarity between the generated image and the subject. Compared to graphical distortion, semantic distortion is more important when the user wants to prevent the unauthorized generation of their images. Overall, MetaCloak achieves the best performance among all the considered baselines.
@@ C.1 Results under Standard Training
We present the results of MetaCloak under the standard training setting in Tab. 5. As we can see, MetaCloak can effectively degrade the personalized generation performance of DreamBooth under the standard training setting. In terms of reference-based semantic metrics, we can see that our method is better in data protection than the previous SOTA in terms of IMS-VGG. Since VGGFaceNet is specially trained on facial datasets and thus more aligned with facial representation, we believe the results on IMS-VGG are more convincing than those on IMS-CLIP. In terms of other metrics, our method is also more effective than the previous SOTA.
@@ C.2 More Visualizations
We visualize the generated images of Dreambooth trained on data perturbed by MetaCloak and other baselines in Fig. 11. As we can see, compared to other baselines, MetaCloak can robustly fool DreamBooth to generate images with low quality and semantic distortion under data transformations. In contrast, other baselines are sensitive to data transformation defenses. In this setting, the generation ability of DreamBooth is retained since the generated images are of high quality. These results demonstrate that MetaCloak is more robust in defense of data transformation.
@@ C.3 More Datasets
To validate the effectiveness of MetaCloak on non-face data, we conduct additional experiments on the DreamBooth-official subject dataset Ruiz et al. [2022], which includes 30 subjects of 15 different classes. Nine out of these subjects are live subjects (dogs and cats), and 21 are objects. Two inference prompts are randomly selected for quality evaluation. As shown in Tab 6, the results demonstrate that our method can still successfully degrade the generation performance on those inanimate subjects.
@@
C.4 More Models
To study the effectiveness of MetaCloak across different architectures, we conduct experiments with three different diffusion models, including Stable Diffusion (SD) v2-1, SD v2-1-base, and SD v1-5. Note that MetaCloak leverages the Stable Diffusion v2-1-base for crafting perturbation. As the visualization in Fig. 5, we can see that MetaCloak can effectively degrade the personalized generation performance of DreamBooth under all the considered architectures, demonstrating the transferability of MetaCloak across architectures.
C.5 Training DreamBooth on Replicate
We test the effectiveness of MetaCloak in the wild by training DreamBooth on the Replicate platform [Replicate, 2023]. The Replicate platform is an online training-as-service platform that allows users to upload their own images and train DreamBooth on them. The generated image of the trained DreamBooth is shown in Fig. 6. As we can see, MetaCloak can effectively degrade the personalized generation performance of DreamBooth under this setting. More visualizations are in the App. C.5. As can be seen, MetaCloak can effectively degrade the personalized generation performance of DreamBooth under both Full-FT and LoRA-FT settings, demonstrating that MetaCloak can seriously threaten Dreambooth’s online training services.
@@ C.5 Training DreamBooth on Replicate
C.6 More Results on Adversarial Purification
@@
    where $x$ is the original image, $Z$ is the reconstructed one, and $\lambda_{TV}$ is the weight on TV term. The first term ensures $Z$ stays close to the original image $x$, and the second term minimizes the total variation which reduces noise and enforces smoothness.
@@
C.7 Effectiveness and Stealthiness under Different Radii
To study the effectiveness of MetaCloak under different radii, we conducted experiments with different radii under the Trans. Training setting. As shown in Tab 8, increasing the radius can effectively improve the effectiveness of MetaCloak. However, when the radius is too large, the stealthiness of injected noise will also be compromised since some specific noise patterns will overwhelm the image content, as shown in Fig. 8. We conclude that the study of how to further improve the stealthiness of MetaCloak under large radii is an important future direction.
C.8 Resiliance under Low Poisoning Ratio
To study the effectiveness of MetaCloak under low poisoning rates, we conduct experiments with different poisoning rates from $\{0\%,25\%,50\%,75\%,100\%\}$ under the two training settings. As shown in Tab 7 and Fig. 10, increasing the poisoning rate can effectively improve the effectiveness of MetaCloak. However, when the poisoning rate is too low, the effectiveness of MetaCloak will be compromised since there is some knowledge leakage. How to effectively protect data under a low poisoning rate is an important future direction.
@@ C.8 Resiliance under Low Poisoning Ratio
C.9 Improving Stealthiness with ReColorAdv
To further improve the stealthiness of our method, we explore replacing the $\ell_{\infty}$-norm constrained attack with ReColorAdv Laidlaw and Feizi [2019], which generates adversarial images by applying a single pixel-wise function $f_{g}$ with parameters $g$ to every color value on the LUV color space. To make the perturbation imperceptible, $f_{g}$ is bounded such that for each pixel $\mathbf{c}_{\mathbf{i}}$, $\|f_{g}(\mathbf{c}_{\mathbf{i}})-\mathbf{c}_{\mathbf{i}}\|_{\infty}\leq\epsilon$, and $\epsilon$ is set as 0.06, 0.08 and 0.10 respectively to observe the method’s performance under different restrictions. To enforce the bound, we optimize the parameters $g$ with PGD, projecting $g$ back to the $\epsilon$ ball after every gradient step. The results are shown in Figure 12. As we can see from the figure, ReColorAdv crafts more uniform and global perturbations, which preserves dependencies between features such as the relationship between light and shadow and the shape boundaries. Unlike the $\ell_{\infty}$-norm attack, the defense effect of ReColorAdv is reflected in the odd color of the generated images instead of collapsed patterns and strips. However, ReColorAdv becomes less effective when the inferring prompt is different from the prompt used in the perturbation crafting process, which is "A photo of sks person" in our case. In conclusion, we show that the ReColorAdv is promising for further improving the stealthiness of our method.
Figure 12: Perturbed images by ReColorAdv and generated images of DreamBooth trained on poisoned data.
C.10 Understanding why MetaCloak works
To understand why MetaCloak works, we visualize the learned noise and conduct wavelet analysis on the noise. We first visualize the original images, perturbed images, and the corresponding noise in Fig. 13. As can be seen from Figure.14, the noise learned by MetaCloak is significantly sharper than the noise generated by the previous method. Thus, the pattern of our noise is less likely to be obscured and turn blurry when encountering defensive measures such as Gaussian transformation and is more likely to fool subject-driven generative models such as Dreambooth. We further plot the distribution of the scales of the learned noise in Figure.15. As can be observed, the scales of the noise are polarized, clustering at both extreme values, and thus more resistant to perturbations.
@@ C.10 Understanding why MetaCloak works
Figure 15: The density of noise scale generated by MetaCloak. We rescale the pixel value of noise to the range $[0,255]$ for visualization.

### v3

@@ 7 Conlusion
This paper proposes MetaCloak as the first work that protects user images from unauthorized subject-driven text-to-image synthesis under data transformation purification. We resolve the limitations of existing works in sub-optimal optimization and fragility to data transformations with a novel meta-learning framework and transformation-robust perturbation crafting process. Extensive experiments demonstrate the effectiveness of MetaCloak. Future directions include establishing mechanism interpretations and designing more efficient perturbations. Acknowledgement. In this work, Yixin Liu, Yutong Dai, and Lichao Sun were supported by the National Science Foundation Grants CRII-2246067.
@@ A.1 Hardware and DreamBooth Training Details
All the experiments are conducted on an Ubuntu 20.04.6 LTS (focal) environment with 503GB RAM, 10 GPUs (NVIDIA® RTX® A5000 24GB), and 32 CPU cores (Intel® Xeon® Silver 4314 CPU @ 2.40GHz). Python 3.10.12 and Pytorch 1.13 are used for all the implementations. For the DreamBooth full trianing mode, we use the 8-bit Adam optimizer [24] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ under bfloat16-mixed precision and enable the xformers for memory-efficient training. For calculating prior loss, we use 200 images generated from Stable Diffusion v2-1-base with the class prompt “a photo of a person”. The weight for prior loss is set to 1. For the instance prompt, we use “a photo of sks person”. During the meta-learning process, we regularly delete the temporary models and store the surrogates back to the CPU to save GPU memory. It takes about 3 GPU hours to craft perturbations for an instance under this strategy.
@@ B.1 Evaluation Metrics
In this section, we describe the evaluation metrics used in our experiments in more detail. For our proposed CLIP-IQAC, we calculate the CLIP score difference between “a good photo of [class]” and “a bad photo of [class]”. For calculating SDS and IMS-VGGNet, we leverage the APIs for face recognition and face embedding extraction in the deep face library [48]. In terms of graphical quality, we found that the commonly used metric, BRISQUE [32] is not a faithful metric when we conduct additional data transformations like Gaussian Filtering. We thus omit this score when presenting results in the main text. For instance, the BRISQUE score of a fully poisoned Dreambooth is better than the clean one, as shown in Tab. 8. Among all the metrics considered, we found that SDS and IMS-VGG are more aligned with our perception of evaluating Dreambooth’s personalized generation performance. The SDS score indicates whether a subject is presented in the generated image, while the IMS-VGGNet score measures the similarity between the generated image and the subject. Compared to graphical distortion, semantic distortion is more important when the user wants to prevent the unauthorized generation of their images. Overall, MetaCloak achieves the best performance among all the considered baselines.
@@ C.1 Results under Standard Training
We present the results of MetaCloak under the standard training setting in Tab. 5. As we can see, MetaCloak can effectively degrade the DreamBooth’s personalized generation performance under the standard training setting. In terms of reference-based semantic metrics, we can see that our method is better in data protection than the previous SOTA in terms of IMS-VGG. Since VGGFaceNet is specially trained on facial datasets and thus more aligned with facial representation, we believe the results on IMS-VGG are more convincing than those on IMS-CLIP. In terms of other metrics, our method is also more effective than the previous SoTA.
@@ C.2 More Visualizations
We visualize the generated images of Dreambooth trained on data perturbed by MetaCloak and other baselines in Fig. 11. As we can see, compared to other baselines, MetaCloak can robustly fool DreamBooth into generating images with low quality and semantic distortion under data transformations. In contrast, other baselines are sensitive to data transformation defenses. In this setting, DreamBooth’s generation ability is retained since the images generated are of high quality. These results demonstrate that MetaCloak is more robust in defense of data transformation.
@@ C.3 More Datasets
To validate the effectiveness of MetaCloak on non-face data, we conduct additional experiments on the DreamBooth-official subject dataset [43], which includes 30 subjects of 15 different classes. Nine out of these subjects are live subjects (dogs and cats), and 21 are objects. Two inference prompts are randomly selected for quality evaluation. As shown in Tab. 6, the results demonstrate that our method can still successfully degrade the generation performance on those inanimate subjects.
@@
C.4 Training DreamBooth on Replicate
We test the effectiveness of MetaCloak in the wild by training DreamBooth on the Replicate platform [40]. The Replicate platform is an online training-as-service platform that allows users to upload their own images and train DreamBooth on them. The generated image of the trained DreamBooth is shown in Fig. 6. As we can see, MetaCloak can effectively degrade DreamBooth’s personalized generation performance in this setting. As can be seen, MetaCloak can effectively degrade the personalized generation performance of DreamBooth under both Full-FT and LoRA-FT settings, demonstrating that MetaCloak can seriously threaten Dreambooth’s online training services.
@@ C.4 Training DreamBooth on Replicate
C.5 More Results on Adversarial Purification
@@
    where $x$ is the original image, $Z$ is the reconstructed one, and $\lambda_{TV}$ is the weight on TV term. The first term ensures $Z$ stays close to the original image $x$, and the second term minimizes the total variation, which reduces noise and enforces smoothness.
@@
More Results against Advanced Purification. To demonstrate the effectiveness of our method against the latest purification, we additionally present the performance of MetaCloak against two more recent purification methods, DiffPure [34] and IMPRESS [19] in Tab. 7. We set the purification iteration for IMPRESS as $N=3$k, and the optimal perturbation time steps for DiffPure as $t^{*}=8$. The results show that these purifications still can’t fully recover the original generation performance, demonstrating MetaCloak’s robustness.
Defenses SDS $IMS_{\text{CLIP}}$ $IMS_{\text{VGG}}$ CLIP-IQAC LIQE
$\times$ 0.401 $\pm$ 0.485 0.641 $\pm$ 0.118 -0.225 $\pm$ 0.852 -0.314 $\pm$ 0.266 0.445 $\pm$ 0.497
+IMPRESS 0.541 $\pm$ 0.493 0.691 $\pm$ 0.080 -0.074 $\pm$ 0.858 -0.307 $\pm$ 0.222 0.789 $\pm$ 0.408
+DiffPure 0.743 $\pm$ 0.430 0.640 $\pm$ 0.081 0.361 $\pm$ 0.673 0.020 $\pm$ 0.289 0.938 $\pm$ 0.242
Oracle* 0.903 0.790 0.435 0.329 0.984
Table 7: Resilience against advanced defenses on VGGFace2 under Trans. Training. Oracle* denotes training on clean data.
C.6 Trade-off of Effectiveness and Stealthiness
To study the effectiveness of MetaCloak under different radii, we conducted experiments with different radii under the Trans. Training setting. As shown in Tab. 9, increasing the radius can effectively improve the effectiveness of MetaCloak. However, when the radius is too large, the stealthiness of injected noise will also be compromised since some specific noise patterns will overwhelm the image content, as shown in Fig. 7. We conclude that the study of how to further improve the stealthiness of MetaCloak under large radii is an important future direction.
C.7 Resilience under Low Poisoning Ratio
To study the effectiveness of MetaCloak under low poisoning rates, we conduct experiments with different poisoning rates from $\{0\%,25\%,50\%,75\%,100\%\}$ under the two training settings. As shown in Tab. 8 and Fig. 10, increasing the poisoning rate can effectively improve the effectiveness of MetaCloak. However, when the poisoning rate is too low, the effectiveness of MetaCloak will be compromised since there is some knowledge leakage. How to effectively protect data under a low poisoning rate is an important future direction.
@@ C.7 Resilience under Low Poisoning Ratio
Figure 12: Effect of each component in MetaCloak on the quality of the generated images. The inferring prompt is “a dslr photo of a person”.
C.8 Improving Stealthiness with ReColorAdv
To further improve the stealthiness of our method, we explore replacing the $\ell_{\infty}$-norm constrained attack with ReColorAdv [25], which generates adversarial images by applying a single pixel-wise function $f_{g}$ with parameters $g$ to every color value on the LUV color space. To make the perturbation imperceptible, $f_{g}$ is bounded such that for each pixel $\mathbf{c}_{\mathbf{i}}$, $\|f_{g}(\mathbf{c}_{\mathbf{i}})-\mathbf{c}_{\mathbf{i}}\|_{\infty}\leq\epsilon$, and $\epsilon$ is set as 0.06, 0.08 and 0.10 respectively to observe the method’s performance under different restrictions. To enforce the bound, we optimize the parameters $g$ with PGD, projecting $g$ back to the $\epsilon$ ball after every gradient step.
The results are shown in Fig. 13. As we can see from the figure, ReColorAdv crafts more uniform and global perturbations, which preserves dependencies between features such as the relationship between light and shadow and the shape boundaries. Unlike the $\ell_{\infty}$-norm attack, the defense effect of ReColorAdv is reflected in the odd color of the generated images instead of collapsed patterns and strips. However, ReColorAdv becomes less effective when the inferring prompt is different from the prompt used in the perturbation crafting process, which is "A photo of sks person" in our case. In conclusion, we show that the ReColorAdv is promising to improve the stealthiness of our method further.
Figure 13: The Perturbed images by ReColorAdv and the generated images of DreamBooth trained on perturbed data.
C.9 Understanding Why MetaCloak works
To understand why MetaCloak works, we visualize the learned noise and conduct wavelet analysis on the noise. We first visualize the original images, perturbed images, and the corresponding noise in Fig. 14. As can be seen from Fig. 15, the noise learned by MetaCloak is significantly sharper than the noise generated by the previous method. Thus, the pattern of our noise is less likely to be obscured and turn blurry when encountering defensive measures such as Gaussian transformation and is more likely to fool subject-driven generative models such as Dreambooth. We further plot the distribution of the scales of the learned noise in Fig. 16. As can be observed, the scales of the noise are polarized, clustering at both extreme values, and thus more resistant to perturbations.
@@ C.9 Understanding Why MetaCloak works
Figure 16: The density of noise scale generated by MetaCloak. We re-scale the pixel value of noise to the range $[0,255]$ for visualization.